OpenStudy (anonymous):

Help please !!!!!! Solve and find out what X and Y equals using subsitution. 3x-y=-8 2x+5y=-45 Explain and Show all your work! Thank you :) 

OpenStudy (jdoe0001):

ok.... so... try first solving the 1st equation for "y", see what you get

OpenStudy (mertsj):

The asker is offline. Not too interested, I'd guess.

OpenStudy (polaris_s0i):

First solve one equation for y \[y = -3x -8\] now substitute y into the other equation: \[2x + 5(-3x - 8) = -45\] \[2x - 15x - 40 = -45\] \[-13x = -5\] \[x = \frac{5}{13}\] Then to validate find y in 2 \[2\frac{5}{13} + 5y = -45\] \[\frac{10}{13} + 5y = -45\] \[5x = \frac{-585}{13} - \frac{10}{13}\] \[x = -\frac{119}{13}\] then check in 1 as well: \[y = -3(\frac{5}{13}) -8\] \[y = \frac{-15}{13} - \frac{104}{13}\] \[y = -\frac{119}{13}\]

OpenStudy (anonymous):

oh, ok Thank You !!!!!!!!!!!!!! I get it now. sorry that i was offline. my laptop went out of battery :) but thanks again for your help polaris_s0i ! XD

OpenStudy (mertsj):

The solution shown is incorrect.

OpenStudy (mertsj):

y=3x+8